OpenStudy (anonymous):

\[f(x)=cosh(x)=1+0+\frac{x^2}{2!}+0+\frac{x^4}{4!}+....\] Maclaurin's I'm having some trouble writing this in summation form \[\sum_{n=0}^{\infty} \frac{f^(0)}{n!}x^(n)\] I know that: All even #'s can be written as 2n All odd #'s can be written as 2n+1 \[\sum_{n=0}^{\infty} \frac{...x^{(2n+1)}}{(2n+1)!}\] I think? so how do I write the \[f^n (0)\] correctly?

OpenStudy (anonymous):

here is my table n f^n x f^n 0 0 coshx 1 1 sinhx 0 2 coshx 1 3 sinhx 0 4 coshx 1

OpenStudy (turingtest):

these are even values of the exponent, so we use the other formula with the 2n

OpenStudy (anonymous):

ooops

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} \frac{...x^{(2n)}}{(2n)!}\]

OpenStudy (turingtest):

besides that you have written all there is to the series; there is nothing in front of each term that you are missing in your sigma notation

OpenStudy (turingtest):

so just take out the dots lol

OpenStudy (anonymous):

oooh because their just zero's and one's ?

OpenStudy (turingtest):

yup try out some values of n plug in n=0 n=1 n=2 etc. you will see you get your series

OpenStudy (anonymous):

check out my table

OpenStudy (turingtest):

you do not need to represent the 0 terms

OpenStudy (turingtest):

you only need to represent the resulting series

OpenStudy (anonymous):

\[f(x)=cosh(x)=1+\frac{x^2}{2!}+\frac{x^4}{4!}+....\] ?

OpenStudy (turingtest):

your table is for the coefficients, and it should tell you that you are keeping only even exponents, hence you use 2n for the exponent and factorial

OpenStudy (turingtest):

that agrees with your results doesn't it?

OpenStudy (anonymous):

makes sense

OpenStudy (turingtest):

so there you have it\[\cosh x=\sum_{n=0}^\infty{x^{2n}\over (2n)!}\]a nice, simple series representation

OpenStudy (anonymous):

\[f(x)=cosh(x)=1+\frac{x^2}{2!}+\frac{x^4}{4!}+....=\sum_{n=0}^{\infty} \frac{x^{(2n)}}{(2n)!}\] final answer. Very good, THankss!

OpenStudy (turingtest):

welcome!